# Data-Efficient Time-Dependent PDE Surrogates: Graph Neural Simulators vs Neural Operators

### v1

@@ 4 Conclusions
In this work, we performed a systematic comparison of generalization performance of various NO frameworks, including DeepONet and FNO-based variants with Graph Neural Simulators (GNS), on three representative two-dimensional PDEs: the viscous Burgers equation, the coupled Burgers system, and the Allen–Cahn equation. Across all cases, we evaluated both qualitative solution contours and quantitative relative $L_{2}$ errors under varying amounts of training data. Our results demonstrate that DeepONet-based methods are more sensitive to limited training data compared to other NOs, with DON-AR consistently exhibiting the largest errors due to the accumulation of approximation errors over time, which is further exacerbated in low-data regimes. TI-based DeepONet variants improve upon the generalization performance of DON-FR, particularly when training data is very limited. However, with sufficient training trajectories, DON-FR surpasses the TI variants due to a more accurate representation of the output solution space, enabled by the increased availability of data. FNO frameworks consistently outperform all DeepONet variants, especially for PDEs whose solutions are dominated by low-frequency modes, making them particularly amenable to spectral-space learning and enabling efficient capture of the dominant frequency modes that govern the PDE dynamics with relatively few training samples.
Notably, GNS emerges as the most robust and accurate framework across all PDEs, even in low-data regimes. Its superior performance is attributed to the explicit modeling of spatial interactions and local dynamics through graph-based node and edge feature representations, allowing it to naturally capture complex, nonlinear, and coupled phenomena such as viscous dissipation, nonlinear cross-advection, and interfacial evolution. This inductive bias enables GNS to learn physically consistent representations, making it particularly effective for long-horizon rollouts and low-data scenarios. Overall, our study highlights the strengths and limitations of current neural operator approaches, demonstrating that while kernel integral operators such as FNO, which operate in spectral space, are highly effective for certain PDEs, graph-based architectures like GNS offer a more universal, physics-aware framework for learning spatiotemporally complex dynamics, particularly in low-data regimes.
@@ Appendix A Computational Costs
Problem Method Batch size Training time (iter/sec) $N_{test}$ Inference time (sec)
2D Burgers scalar GNS 2 72.87 1000 181.535
@@ B.1 Graph Neural Simulators (GNS)
2D Allen-Cahn Node: [7,64,64] Edge: [7,64,64] 6 MP layers [64,64,1] GELU 500 4
@@ B.2 Fourier Neural Operator (FNO)-based variants
Table A3: FNO architecture and training settings across different PDE examples. In all the cases, 4 Fourier layer blocks are used.

### v2

@@ 4 Conclusions
In this work, we performed a systematic comparison of generalization performance of various NO frameworks, including DeepONet and FNO-based variants with Graph Neural Simulators (GNS), on four representative two-dimensional PDEs: the viscous Burgers equation, the coupled Burgers system, the Allen–Cahn equation, and the nonlinear shallow water equations. Across all cases, we evaluated both qualitative solution contours and quantitative relative $L_{2}$ errors under varying amounts of training data. Our results demonstrate that DeepONet-based methods are more sensitive to limited training data compared to other NOs, with DON-AR consistently exhibiting the largest errors due to the accumulation of approximation errors over time, which is further exacerbated in low-data regimes. TI-based DeepONet variants improve upon the generalization performance of DON-FR, particularly when training data is very limited. However, with sufficient training trajectories, DON-FR surpasses the TI variants due to a more accurate representation of the output solution space, enabled by the increased availability of data. FNO frameworks consistently outperform all DeepONet variants, especially for PDEs whose solutions are dominated by low-frequency modes, making them particularly amenable to spectral-space learning and enabling efficient capture of the dominant frequency modes that govern the PDE dynamics with relatively few training samples.
Notably, GNS emerges as the most robust and accurate framework across all PDEs, even in low-data regimes. Its superior performance is attributed to the explicit modeling of spatial interactions and local dynamics through graph-based node and edge feature representations, allowing it to naturally capture complex, nonlinear, and coupled phenomena such as viscous dissipation, nonlinear cross-advection, interfacial evolution, and wave-height coupling with gravity-driven free-surface flow. This inductive bias enables GNS to learn physically consistent representations, making it particularly effective for long-horizon rollouts and low-data scenarios. Overall, our study highlights the strengths and limitations of current neural operator approaches, demonstrating that while kernel integral operators such as FNO, which operate in spectral space, are highly effective for certain PDEs, graph-based architectures like GNS offer a more universal, physics-aware framework for learning spatiotemporally complex dynamics, particularly in low-data regimes.
@@ Appendix A Computational Costs
Problem Method Batch size
Training time
(iter/sec)
$N_{test}$
Inference time
2D Burgers scalar GNS 4 72.87 1000 181.535
@@ B.1 Graph Neural Simulators (GNS)
2D Allen Cahn Node: [7,64,64] Edge: [7,64,64] 6 MP layers [64,64,1] GELU 500 4
2D nonlinear SWE Node: [9,64,64] Edge: [9,64,64] 6 MP layers [64, 64, 3] GELU 600 2
@@ B.2 Fourier Neural Operator (FNO)-based variants
Table A3: FNO architecture and training settings across different PDE examples. Except for the 2D nonlinear SWE case, where 6 Fourier blocks are used, all other cases use 4 Fourier blocks.
@@ Appendix C Additional Results
C.4 2D nonlinear Shallow Water Equations
Figure A5: 2D nonlinear shallow water equations: Predicted solution contours of all frameworks on 50 training trajectories, evaluated on 500 trajectories for predicting the $u$-velocity field of a representative sample.
Figure A6: 2D nonlinear shallow water equations: Predicted solution contours of all frameworks on 50 training trajectories, evaluated on 500 trajectories for predicting the $v$-velocity field of a representative sample.
Figure A7: 2D nonlinear shallow water equations: Predicted solution contours of all frameworks on 50 training trajectories, evaluated on 500 trajectories for predicting the height field, $h$, of a representative sample.